OpenStudy (anonymous):

The product of a number and 2 is 4 less than the quotient of a number and 3. Use x as a variable. Need help writing this into an equation/expression!

OpenStudy (texaschic101):

2x = (x/3) - 4

OpenStudy (texaschic101):

product is a result of multiplication...quotient is the result of division

OpenStudy (anonymous):

i got 2x-4=x/3. how come the -4 is on the right side, not the left? is it because "is" means an equal sign?

OpenStudy (texaschic101):

is means equals

OpenStudy (anonymous):

thanks you so much for the clarification! helped me a lot

OpenStudy (texaschic101):

sure thing :) this can be easily solved by multiplying the entire equation by 3. Then you s=do not have to mess with fractions

OpenStudy (texaschic101):

oops...typo...then you do not have to mess with fractions

OpenStudy (anonymous):

thank you :D i dont have to solve it but that's a good reminder :) have a great night thank u so much :D

OpenStudy (texaschic101):

you too....good night :)